OpenStudy (anonymous):

find the product of (3x-7)²

OpenStudy (anonymous):

look at the equation as (3x-7)(3x-7)

OpenStudy (anonymous):

means????? 9x²+21x-21x-45?? am i right?

OpenStudy (anonymous):

\[(3x-7) ^2 \] \[(3x^2)-49\] \[x^2=49-9\] \[x^2 = 40 \] \[x =6.324\] =)

OpenStudy (anonymous):

ty hahah